OpenStudy (anonymous):

Determine whether the function y=x^2 is a solution of the differential equation 2xy' + 4y=10x^3

OpenStudy (anonymous):

can i say yes it is ?

OpenStudy (anonymous):

yes is the answer, but how did you get it? o.o

OpenStudy (anonymous):

im still learning but if you differentiate the problem

OpenStudy (anonymous):

you end up with 2y'+4=10x^2 6 =10 x^2 6/10 = x^2 3/5= x^2 i dunno is that how you get the answer ?

OpenStudy (anonymous):

to be honest i'm not really sure what you did XD but I think i have to do something with integrals

OpenStudy (anonymous):

i used the power rule :P or tried

OpenStudy (anonymous):

then i used my logic super wacked (logic) and got the answer

OpenStudy (anonymous):

then i will probably be missing somthing of course

OpenStudy (ash2326):

y=x^2 y'=2x multiply 2x 2xy'=4x^2 y=x^2 2xy'=4y=4x^2 2xy'+4y=2*4x^2 so 2xy'+4y=8x^2 so it is not the solution of the differential equation

OpenStudy (anonymous):

2xy' + 4y=10x^3 y= x^2 y'= 2x Plug in those values: 2x(2x) + 4(x^2)= 10x^3 Solve: 4x^2 + 4x^2= 10x^3 LHS DOES NOT = RHS THEREFORE NOT A SOLUTION